# Bound Orbits and Gravitational Wave Radiation Around the Hairy Black Hole

### v2

@@ I Introduction
In this paper, we mainly discuss the periodic orbits of a timelike test particle in the background of a regular hairy black hole and preliminarily study the characteristics of gravitational waves generated by periodic orbits in the EMRI system in this background. The structure of this paper is as follows: In Section 2, we give a brief introduction to the regular hairy black hole and calculate the effective potential of a timelike test particle around it. In Section 3, in the first subsection, we introduce the relevant properties of MBO and ISCO; in the second subsection, we introduce precession and periodic orbits. In Section 4, we conduct a preliminary exploration of the gravitational waves radiated by periodic orbits. Finally, in Section 5, a summary is given. In this paper, unless otherwise specified, the natural unit system with $c=G=1$ is used throughout.
@@ II The regular hairy black hole
In this paper, we briefly review the black hole solutions obtained by Jorge Ovalle and other scholars using the gravitational decoupling analysis method. In their calculations, they introduced the tensor vacuum $\theta_{\mu\nu}$, an additional energy-momentum tensor used to describe a new matter field. By applying geometric deformations to the traditional Schwarzschild solution, this approach effectively avoids the central singularity problem. Under the conditions of satisfying the weak energy condition and the event horizon constraint, the authors ultimately derived a regular hairy black hole solution, with its metric given as followsOvalle et al. (2023)
@@
Here, $M$ represents the ADM mass of the black hole, and $\alpha$ is the hair parameter, which can be interpreted as the source of the deformation of the Minkowski vacuum. As stated in the literatureOvalle et al. (2023), when $\alpha\to 0$, the metric reduces to Minkowski spacetime, indicating the absence of any gravitational field. When $\alpha\to\infty$, the metric converges to the Schwarzschild solution in the static case, exhibiting gravitational characteristics similar to those of a classical black hole. In addition, this solution possesses a regular spacetime structure without a central singularity, thereby overcoming the singularity problem of the traditional Schwarzschild black hole.

### v3

@@ I Introduction
In this paper, we mainly discuss the periodic orbits of a timelike test particle in the background of a regular hairy black hole and preliminarily study the characteristics of gravitational waves generated by periodic orbits in the EMRI system in this background. The structure of this paper is as follows: In Section II, we give a brief introduction to the regular hairy black hole and calculate the effective potential of a timelike test particle around it. In Section III, in the first subsection, we introduce the relevant properties of MBO and ISCO; in the second subsection, we introduce precession and periodic orbits. In Section IV, we conduct a preliminary exploration of the gravitational waves radiated by periodic orbits. Finally, in Section V, a summary is given. In this paper, unless otherwise specified, the natural unit system with $c=G=1$ is used throughout.
@@ II The regular hairy black hole
In this paper, we briefly review the black hole solutions obtained by Jorge Ovalle and other scholars using the gravitational decoupling analysis method. In their calculations, they introduced the tensor vacuum $\theta_{\mu\nu}$, an additional energy-momentum tensor used to describe a new matter field. By applying geometric deformations to the traditional Schwarzschild solution, this approach effectively avoids the central singularity problem (The necessary formula derivations are provided in Appendix.). Under the conditions of satisfying the weak energy condition and the event horizon constraint, the authors ultimately derived a regular hairy black hole solution, with its metric given as followsOvalle et al. (2023)
@@
Here, $M$ represents the ADM mass of the black hole, and the parameter $\alpha$ is the hair parameter, which represents the source that causes the deformation of the Minkowski vacuum. As $\alpha\to 0$, the metric degenerates into the flat Minkowski spacetime, indicating that there is no gravitational field and the spacetime returns to a flat structure without gravity. On the other hand, as $\alpha\to\infty$, the metric converges to the static Schwarzschild solution, exhibiting gravitational characteristics similar to those of a classical black hole. In this process, the variation of $\alpha$ causes the spacetime geometry to first transition from flat spacetime to a regular hairy black hole solution with a hair structure, and further transition to the classical Schwarzschild black hole solution. Therefore, the existence of the regular hairy black hole is closely related to the value of $\alpha$, indicating that this parameter plays a decisive role in the evolution of the gravitational field and spacetime structure. Moreover, this solution has a regular spacetime structure and does not have a central singularity, thus overcoming the singularity problem of the traditional Schwarzschild black hole.
@@ Acknowledgements
Appendix
In this appendix, we will briefly outline the derivation process of the metric (2), explaining how the metric is derived from the Einstein-Hilbert action through the gravitational decoupling method and satisfies the Einstein field equations, while also providing the corresponding Lagrangian density.
Jorge Ovalle and other scholars adopted the gravitational decoupling method, starting from the Einstein-Hilbert action, which is expressed as followsOvalle et al. (2023)
$S=\int{\left[\frac{R}{2\kappa}+\mathcal{L}_{M}+\mathcal{L}_{\Theta}\right]\sqrt{-g}}d^{4}x,$ (35)
here, $R$ is the Ricci scalar, $\mathcal{L}_{M}$ is the Lagrangian density containing the standard matter, and $\mathcal{L}_{\Theta}$ is the Lagrangian density introduced for additional sources or gravitational sectors beyond general relativity.
For these two sources, their energy-momentum tensors can be written as
$T_{\mu\nu}=-\frac{2}{\sqrt{-g}}\frac{\delta\left(\sqrt{-g}\mathcal{L}_{M}\right)}{\delta g^{\mu\nu}}=-2\frac{\delta\mathcal{L}_{M}}{\delta g^{\mu\nu}}+g_{\mu\nu\mathcal{L}_{M}},$ (36)
$\theta_{\mu\nu}=-\frac{2}{\sqrt{-g}}\frac{\delta\left(\sqrt{-g}\mathcal{L}_{\Theta}\right)}{\delta g^{\mu\nu}}=-2\frac{\delta\mathcal{L}_{\Theta}}{\delta g^{\mu\nu}}+g_{\mu\nu\mathcal{L}_{\Theta}}.$ (37)
At this point, the Einstein field equations can be written as
$G_{\mu\nu}\equiv R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}=\kappa\left(T_{\mu\nu}+\theta_{\mu\nu}\right),$ (38)
here $\kappa=8\pi G$.
For a static spherically symmetric solution, it can typically be expressed as
$ds^{2}=-e^{\nu\left(r\right)}dt^{2}+e^{\lambda\left(r\right)}dr^{2}+r^{2}d\Omega^{2}.$ (39)
Substituting the above expression into the field equation (38) yields the corresponding energy-momentum tensor
$T_{\mu}^{\nu}=\mathrm{diag}\left[-\varepsilon,p_{r},p_{\theta},p_{\theta}\right],$ (40)
$\theta_{\mu}^{\nu}=\mathrm{diag}\left[-\mathcal{E},\mathcal{P}_{r},\mathcal{P}_{\theta},\mathcal{P}_{\theta}\right].$ (41)
The detailed derivation process can be found in the original referenceOvalle et al. (2023).
In order to obtain the regular hairy black hole solution, we use the traditional black hole solution as a seed source, and on this basis, introduce additional matter sources to avoid the emergence of singularities. The expression for the seed source is
$ds^{2}=-e^{\xi\left(r\right)}dt^{2}+e^{\mu\left(r\right)}dr^{2}+r^{2}d\Omega^{2}.$ (42)
After adding the additional source $\theta_{\mu\nu}$, the corresponding function can be written as
$\xi\rightarrow\nu=\ \xi+\ \alpha g,$ (43)
$e^{-\mu}\rightarrow e^{-\lambda}=e^{-\mu}+\alpha f.$ (44)
Here, $\alpha$ is the deformation parameter, and $g$ and $f$ are correction functions related to the matter source.
Combining expressions (43) and (44), the Einstein equation (38) can be split into two sets. The first set is given by $T_{\mu\nu}$, and at this point, we have
$\begin{split}&\kappa\varepsilon=\frac{1}{r^{2}}-e^{-\mu}\left(\frac{1}{r^{2}}-\frac{\mu^{\prime}}{r}\right),\\ &\kappa p_{r}=-\frac{1}{r^{2}}+e^{-\mu}\left(\frac{1}{r^{2}}+\frac{\xi^{\prime}}{r}\right),\\ &\kappa p_{\theta}=\frac{e^{-\mu}}{4}\left(2\xi^{\prime\prime}+\xi^{\prime 2}-\mu^{\prime}\xi^{\prime}+2\frac{\xi^{\prime}-\mu^{\prime}}{r}\right).\end{split}$ (45)
The second set is given by the source $\theta_{\mu\nu}$, and at this point, we have
$\begin{split}&\kappa\mathcal{E}=-\frac{\alpha f}{r^{2}}-\frac{\alpha f^{\prime}}{r},\\ &\kappa\mathcal{P}_{r}-\alpha Z_{1}=\alpha f\left(\frac{1}{r^{2}}+\frac{\nu^{\prime}}{r}\right),\\ &\kappa\mathcal{P}_{\theta}-\alpha Z_{2}=\frac{\alpha f}{4}\left(2\nu^{\prime\prime}+\nu^{\prime 2}+2\frac{\nu^{\prime}}{r}\right)-\frac{\alpha f^{\prime}}{4}\left(\nu^{\prime}+\frac{2}{r}\right).\end{split}$ (46)
where
$Z_{1}=\frac{e^{-\mu}g^{\prime}}{r},$ (47)
$4Z_{2}=e^{-\mu}\left(2g^{\prime\prime}+\alpha g^{\prime 2}+\frac{2g^{\prime}}{r}+2g^{\prime}\xi^{\prime}-\mu^{\prime}g^{\prime}\right).$ (48)
The energy exchange between these two sources is given by the conservation equation
$\nabla_{\sigma}T_{\nu}^{\sigma}=-\frac{\alpha g^{\prime}}{2}\left(\varepsilon+p_{r}\right)\delta_{\nu}^{\sigma}=-\nabla_{\sigma}\theta_{\nu}^{\sigma}.$ (49)
In order to construct a regular hairy black hole solution within the framework of general relativity, the authors of the original paper use the Schwarzschild black hole solution ($T_{\mu\nu}$) as a seed source and prevent the formation of singularities due to gravitational collapse by introducing an additional matter source $\theta_{\mu\nu}$, such as the introduction of a tensor vacuum source. On this basis, they further require that the solution has a well-defined event horizon structure and satisfies the weak energy condition.
By applying these constraint conditions, the regular hairy black hole solution is ultimately obtainedOvalle et al. (2023)
$e^{\nu}=e^{-\lambda}=1-\frac{2M}{r}+\frac{e^{-\alpha r/M}}{rM}\left(\alpha^{2}r^{2}+2M\alpha r+2M^{2}\right).$ (50)
This metric describes the regular black hole solution after the deformation is introduced through gravitational decoupling and the addition of sources. The hair parameter $\alpha$ controls the structure of the metric. As $\alpha\to 0$, the metric degenerates into Minkowski spacetime, and as $\alpha\to\infty$, the metric recovers the standard Schwarzschild solution.
In addition, the metric (50) can also be derived through the variational principle from the assumed action. The total Lagrangian density $\mathcal{L}$ (which includes the additional matter source) can be obtained using the P-dual formalismAyon-Beato and Garcia (1998). In this approach, the expression for the total action is
$S=\int{\left[\frac{R}{2\kappa}+\mathcal{L}\right]\sqrt{-g}}d^{4}x.$ (51)
The specific form of the Lagrangian density $\mathcal{L}$ isOvalle et al. (2023)
$\mathcal{L}\left(P\right)=\frac{\alpha^{3}}{2\kappa{M}^{3}}\left[\xi\left(P\right)-2{M}\right]e^{-\xi\left(P\right)/{M}},$ (52)
here
$\xi\left(P\right)=\alpha\left(\frac{-2\alpha^{2}}{\kappa^{2}P}\right)^{\frac{1}{4}}.$ (53)
The detailed derivation of this part can be found in the reference Ovalle et al. (2023).